338746721494:

Part A: Solve the equation 4(x + 5) = 9x + 4x − 34, and show your work. (5 points) Part B: Use this list of properties to justify each step of your math work. Justifications may be used more than once if needed. (5 points)

Vocaloid:

missing list for part B) Distributive Property Subtraction Property of Equality Combine Like Terms Multiplication Property of Equality Division Property of Equality Given Addition Property of Equality

Vocaloid:

for proofs, you typically start with "given" and copy down the information you are already given so you would start with: Given: 4(x + 5) = 9x + 4x − 34 from there, you can start performing operations, justifying each step with one of the properties from the list. for example, you can distribute the 4 from 4(x+5) using the distributive property of equality. try solving it, I'll help you out if you get stuck.

annacam:

so basically it would be x+6 right

Vocaloid:

yes, that's the solution, but keep in mind the problem is asking for all steps

annacam:

oh so i would have to put how i got it

Vocaloid:

yes

annacam:

oh

Mathyg:

4(x + 5) = 9x + 4x − 34 4x+20=13x-34 20+34=13x-4x 54=9x 6=x